OpenStudy (anonymous):

A football is kicked at ground level with a speed of 18.0 m/s at an angle of 35.0 degrees to the horizontal. How much later does it hit the ground?

OpenStudy (anonymous):

 x- direction v = 18 cos 35 = 14.745 m/s y- direction initial velocity = u = 18 sin 35 = 10.324m/s a = 9.8 m/s^2 final velocity = v = 0 \[t = (v-u) \div a\] t = (0 - 10.324) / -9.8 = 1.05s Since this is the time it takes to just go up, we double it to know the total time. Hence, t = 1.05 X 2 = 2.11s It hits the ground 2.11s later.